OpenStudy (anonymous):

Verify identity for sec^4x -sec^2x= 1/cot^4x + 1/cot^2x

OpenStudy (anonymous):

sec^2x(sec^x-1) = (1+tan^2x)(tan^2x) = tan^2x+tan^4x = 1/cot^2x+1/cot^4x (proved)

OpenStudy (anonymous):

R.H.S=L.H.S